OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - 1/2y What is the value of x? 9/7 24/14 33/14

terenzreignz (terenzreignz):

You want the answer now, or do you want the steps?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

cuz i have a short time to answer 25 questions

terenzreignz (terenzreignz):

it's 33/14

OpenStudy (anonymous):

an not all r easy thx a lot

terenzreignz (terenzreignz):

Just for reference: I took the second equation x = 3 - (1/2)y played with it a bit: (1/2)y = 3 - x y = 2(3 - x) y = 6 - 2x Now that I have a value for y in terms of x, I substitute it for the y in the first equation 4x - 5y = 3 4x - 5(6 - 2x) = 3 4x - 30 + 10x = 3 14x - 30 = 3 14x = 33 x = 33/14 END

OpenStudy (anonymous):

dang u make it look easy

terenzreignz (terenzreignz):

Practice makes perfect :)

OpenStudy (anonymous):

ya thx